# Information-theoretic description of a feedback-control Kuramoto model

### v1

@@ Abstract
Semantic Information Theory (SIT) offers a new approach to evaluating the information architecture of complex systems. In this study we describe the steps required to operationalize SIT via its application to dynamical problems. Our road map has four steps: (1) separating the dynamical system into agent-environment sub-systems; (2) choosing an appropriate coarse graining and quantifying correlations; (3) identifying a measure of viability; (4) implementing a scrambling protocol and measuring the semantic content. We apply the road map to a model inspired by the neural dynamics of epileptic seizures whereby an agent (a control process) attempts to maintain an environment (a base process) in a desynchronized state. The synchronization dynamics is studied through the well-known Kuramoto model of phase synchronization. Our application of SIT to this problem reveals new features of both semantic information and the Kuramoto model. For the latter we find articulating the correlational structure for agent and environment(the oscillators), allows us to cast the model in in a novel computational (information theoretic) perspective, where the agent-environment dynamics can be thought of as analyzing a communication channel. For the former we find that all the information in our system is semantic. This is in contrast to previous SIT studies of foragers in which semantic thresholds where seen above which no further semantic content was obtained.
@@ I Introduction
While extensive discussions of, and definitions for, semantic information exist, they have mostly focused on issues in philosophy or applications within computer science D Polani [2001]; Thompson and Stapleton [2009]; Nehaniv et al. [2002]; Barham [1996]; Deacon [2007]; Corning [2007]; Gleiser and Sowinski [2018]. What has been lacking for a physics of life is a mathematically precise and scientific applicable theory of semantic information that can inform a wide range of domains ranging from neuroscience, collective microbial behavior, social physics and Origin of Life studies Acebrón et al. [2005]; García-Selfa et al. [2021]; Cooper et al. [2013]; Mimar et al. [2019]; COHEN [1977]; Rooney et al. [2006]; Xavier et al. [2020]; Blain and Szostak [2014]; Mimar et al. [2021].
@@
The KW18 formalism appears to offer an attractive path forward for the development of a working theory of semantic information. The difficulty, however, lies in the computational cost of applying the formalism to real world applications. Calculating the information theoretic measures, require a precise elucidation of the state spaces, and the computation of the joint probability distributions between the agent and the environment. Unfortunately such state spaces are often of such high dimensionality that the formalism is comptuationally intractable. Thus if SIT is to live up to its promise, work must be done to find effective measures for the quantities required in its formalism. This includes best practices for definitions of viability, proper mechanisms for scrambling shared information in the intervened tracjectories and, finally, practical methods of dimensional reduction (i.e. coarse-graining) for calculations of information theoretic metrics.
@@
Recently, an attempt to develop such effective versions of the KW18 formalism was presented in Sowinski et al. [2023]. The authors studied a forager model, a minimal model that captures attributes of living systems, namely exploration and resource consumption Bénichou and Redner [2014]; Bhat et al. [2017]; Bénichou et al. [2016]; Bhat and Redner [2022]. The viabilty was defined as the expected lifetime of the forager and the scrambling was done by noising a sensor that the forager used to detect resources. They applied and extended the work of KW18 and demonstrated how semantic information measures provide new insights into forager dynamics. More importantly, they provided a first cut at developing effective measures for application of the SIT formalism to real problems.
@@
An important result from Sowinski et al. [2023] was the identification of a plateau in the viability curve as a function of scrambling correlations between agent and environment. This plateau occurs at at viability threshold that is well below channel capacity. In other words a subset of correlations between agent and environment has no effect on the forager’s lifetime (the information was not semantic). Once the threshold is breached the ability for the forager to survive decreases monotonically with increased scrambling (noise in the sensor). Thus applying the ideas of SIT results in an identification of features in the viability curve that are both novel and independent of the particular foraging strategy used. Whether such semantic saturation is a generic feature of living systems is far from settled, but the results make it clear that finding ways to apply the fundamental principle of SIT results in novel ways to analyze systems from an informational perspective.
@@
In this paper we report on the next step in efforts to operationalize SIT and explore effective versions of the theory. To carry out that task we explore the biophysically motivated problem of a control process (CP) acting on a base process (BP) to maintain overall system viability. We analyze a model of a network of coupled oscillators which tend towards synchronization (the base process) and an intervening agent (the control process) which uses the information on the network to limit synchronization. The synchronization dynamics is studied through the well-known Kuramoto model of phase synchronization Kuramoto [1975]. This setting is loosely inspired by the neurodynamics of epileptic attacks Zheng et al. [2014]; Uhlhaas and Singer [2006]; Siegel et al. [2012]; Ghoshal et al. [2016] where brain pathologies occur due to large scale synchronization of neurons Penfield and Jasper [1954]. Additionally, in versions of the Kuramoto model with non-local coupling Abrams and Strogatz [2004] or symmetry-breaking Zhang and Motter [2021] there exist chimera states, where identically coupled oscillators coexist in synchronized and desysnchronized states. Such states have been observed in both brain networks and the resuscitation of cardiac cells Omel’chenko [2018]. Thus desynchronization is important in many biological contexts. Consequently, the intervening agent here acts as preventing the system from reaching (a potentially) pathological synchronized state. Our goal in this paper, as in Sowinski et al. [2023], is to both explore the application of the KW18 formalism and use SIT to shed new light on a previously well-studied problem, phase-(de)synchronization.
@@
In what is to follow, in Sec. II we explicitly unpack the general steps required for the use of the SIT formalism. In Sec. III.1 we introduce the feedback-control model on the coupled oscillators. We demonstrate that introducing the agent leads to a stochastic version of the original Kuramoto model, with oscillators coupled to a heat-bath mediated by their degree $k_{i}$ and a deformed distribution of frequencies $g(\omega)$. Following this in Sec. III.2 we recast the model in information theoretic terms in which we show that the base process is a stochastic computation, and the control process is akin to analysing a communication channel between the agent and its environment. The viability function is defined in Sec. III.3 and in Sec. III.4 we analyze the semantic content of the correlations between agent/environment demonstrating the absence of a semantic threshold, indicating that every bit of information in the communication channel is crucial to the viability of the system. Finally in Sec. IV we end with a discussion of the implications of our findings. Three Appendixes clarify technical aspects of our model and the approximations used.
@@ II Preliminaries
  • Dynamical System Decomposition. Construct a dynamical system $(S,\mathcal{D})$, where $S$ is the system state space and $D:S\times T\rightarrow S$ is the dynamics on the state space for all times $t\in T$. The dynamics can be deterministic, stochastic, or a combination of the two; if they are continuous then $T\simeq\mathds{R}$ or some subinterval of the reals, and if they are discrete then $T\simeq\mathds{Z}$, or some finite subsequence of the integers. The system must be factorable into an agent and environment, $S=A\times E$, meaning that any state of the system can be written $s=(a,e)$ where $a=(a_{1},a_{2},\dots)\in A$ and $e=(e_{1},e_{2},\dots)\in E$ are, respectively, the agent and environment dynamical degrees of freedom (DoF). We write $s(t)=(e(t),a(t))$ to represent the system evolution through state space, and refer to the specific form of $\mathcal{D}$ responsible for $a(t)$ as the agent’s behavior. $S$ needs to accommodate a (not necessarily unique) ground state for the agent which acts as an attractor for the system dynamics and represents a dead agent and an evolving agent-free environment.
@@
  • Coarse Grain and Quantify Correlations. Correlations between agent and environment DoF are manifest in the information architecture of the system, as quantified by a corpus of measures including mutual information, $I(A:E)$ Cover and Thomas [2006], cross entropy $H(A:E)$ Chen et al. [2022], and transfer entropy $\mathcal{T}_{E\rightarrow A}$ Schreiber [2000]. In practice these quantities may not be calculable or efficiently computable in full, so a bit of finesse is required at this step. A combination of prior knowledge and intuition must be used to coarse grain the system into DoF that lend themselves to the desired analytical or computational techniques, while capturing the minimal dynamics of $S$ required to describe an agent attempting to survive in and coevolve with its environment. In particular, the agent-free attractor must be preserved by the coarse graining. For example, it would be intractable to describe a bacterium’s environment by listing the position and momentum of every atom surrounding it; a simpler description using a coarse-grained glucose density field might suffice. Describing the full biochemistry of the bacterium using every molecular DoF is intractable, but a coarse grained smaller set of DoF such as center of mass position, orientation, stored ATP density, and locomotive modality may not be. Note how this coarse-graining preserves the ground state: in the limit when the bacterium’s stored ATP supply vanishes it is effectively dead.
@@
  • Identify Viability Measure $\mathcal{V}$. Since the ground state remains an attractor for the system dynamics under coarse graining, the distance to it edogenously defines the viability of an agent. Distance here can mean several things: If the attractor is stable, then an agent trajectory will eventually end up on it, i.e. the agent will die. In this case the expected lifetime of the agent is a proper measure of distance, and hence viability. The example of the bacterium falls in this case; the expected first crossing time to reach zero ATP supply is a viability measure. If the attractor is unstable, then the coarse graining from the prior step must be able to identify an order parameter representing the efficacy of the agent at maintaining an out-of-(agent-free)equilibrium environment. The efficacy vanishes in an agent-free environment, and takes some non-zero value once the agent is introduced. The expected efficacy is then a proper measure of distance, and hence viability. The viability of the agent measured in either case is referred to as the actual viability.
@@
  • Scrambling Protocol and measuring semantic Content. The goal of SIT is to identify how the correlations and viability defined in the last two steps relate to one another. Specifically, to find which correlations matter to the agent. This is done by implementing a scrambling protocol on the correlations, and examining the resultant change in viability due to the impact the scrambled correlations have on the dynamics of the system. This was expounded on in KW18 using interventions to generate counterfactual histories, and requires a search over all possible partitions of the system degrees of freedom. Though rigorously defined, the procedure becomes unmanageable for systems with more than several hundred states. Rather than going through all partitions, the coarse graining, if done correctly, identifies which coarse-grained DoF should be investigated and scrambled. In the example mentioned earlier, the bacterium’s orientation and whether it is tumbling or rolling are correlated to gradients and magnitudes of the glucose density. A scrambling protocol could be introduced at the level of the dynamics by intervening on the bacterium’s ability to sense either a magnitude or a gradient, or on the conditions that determine when the bacterium should switch it’s locomotive modality. The protocol should be tuneable, in the sense that the actual viability is a limiting case. To contrast with the actual viability, the scrambled viability is measured as the protocol is tuned higher. The point at which the scrambled viability diverges from the actual viability is the semantic threshhold Sowinski et al. [2023].
@@ III Kuramoto Model with Feedback-Control
Kuramoto’s famous coupled oscillator model has been well-studied since it’s introduction nearly half a century ago Kuramoto [1975]. From the synchronization of fireflies, cardiac pacemakers, laser, power transmission relay networks among others, it has proven a rich starting point for modelling systems in which synchronization is prominent Mirollo and Strogatz [1990]; Strogatz et al. [2005]; Moiseff and Copeland [2010]; Li and Wong [2017]. Above a critical coupling strength, an extensive fraction of oscillators achieve an equilibrium synchronized state, while the desynchronized states are non-equilibrium transients in the dynamics. As mentioned in the introduction, the phenomenon of desynchronization has biological relevance in terms of pathological states in the human brain and in cardiac arrythmia. Inpired by this, we develop a feedback-control model of an agent acting on a population of coupled oscillators that attempt to prevent synchronization. We then apply the SIT framework to this idealized model and develop a new information theoretic perspective on coupled oscillators. In our formulation, the control process (an agent) seeks to drive the base process (the Kuramoto model) to a desired far-from-equilibrium set of states. Remaining far-from-equilibrium provides a definition of viability i.e. it motivates the teleological dynamics of the agent (the control process) seeking to keep the system from complete synchronization. Viewing the system through the lens of semantic information the agent can be interpreted as a repair process acting on a malfunctioning background.
Figure 1: Oscillators $\{\theta_{i}\}$ are placed on the nodes of both power law (upper left) and Poisson graphs (upper right). Kuramoto dynamics are imposed resulting in Kuramoto environments. At each node (lower left) there is a local order parameter and local mean field, $\{r_{i},\phi_{i}\}$. Together, these dynamical quantities define the local mean field coupling for each oscillator of natural frequency $\omega_{i}$. Control agents are placed in the Kuramoto environments, moving uniformly at random along edges, and perturbing the nodes based off of local mean field estimates $\phi_{i}$. Here $\alpha$ describes the strength of the perturbation that the agent applies, and $\beta$ the time-scale or speed of the perturbation.
@@ III.1.1 Environment Description
The environment is a population of $N$ oscillators $(\theta_{1}\dots\theta_{N})$, $\forall i,\theta_{i}\in(-\pi,\pi]$ placed on the nodes of an undirected graph $G$ with connections parameterized by the adjacency matrix ${{\boldsymbol{A}}}$ whose elements are $A_{ij}=1$ if nodes $i$ and $j$ are connected, and 00 otherwise. In what is to follow, we consider Poisson and scale-free degree distributions generated via the configuration model Newman [2018]. The oscillators are coupled to one another with a coupling $\sigma$, which we take to be a function of time. Each oscillator has a natural frequency, $\omega_{i}$, drawn from a distribution $g(\omega)$. Without loss of generality we set $\langle\omega\rangle=0$. The variance, or more precisely the root mean square frequency $\omega_{\text{rms}}=\langle\omega^{2}\rangle^{1/2}$, is used to rescale time, $\tilde{t}=\omega_{\text{rms}}t$. Dropping the tildes, the dynamics of the dimensionalized environment are the Kuramoto equations of motion, thus,
@@
with $g(\omega)$ having vanishing mean and unit variance.

### v3

@@ Abstract
Semantic Information Theory (SIT) offers a new approach to evaluating the information architecture of complex systems. In this study we describe the steps required to operationalize SIT via its application to dynamical problems. Our road map has four steps: (1) separating the dynamical system into agent-environment sub-systems; (2) choosing an appropriate coarse graining and quantifying correlations; (3) identifying a measure of viability; (4) implementing a scrambling protocol and measuring the semantic content. We apply the road map to a model inspired by the neural dynamics of epileptic seizures whereby an agent (a control process) attempts to maintain an environment (a base process) in a desynchronized state. The synchronization dynamics is studied through the well-known Kuramoto model of phase synchronization. Our application of SIT to this problem reveals new features of both semantic information and the Kuramoto model. For the latter we find articulating the correlational structure for agent and environment (the oscillators), allows us to cast the model from a computational (information theoretic) perspective, where the agent-environment dynamics form a communication channel. For the former we find a topological dependence of whether information can be considered semantic or not.
@@ I Introduction
While extensive discussions of, and definitions for, semantic information exist, they have mostly focused on issues in philosophy or applications within computer science D Polani [2001]; Thompson and Stapleton [2009]; Nehaniv et al. [2002]; Barham [1996]; Deacon [2007]; Corning [2007]; Gleiser and Sowinski [2018, 2018]; Sowinski [2016]. What has been lacking for a physics of life is a mathematically precise and scientific applicable theory of semantic information that can inform a wide range of domains ranging from neuroscience, collective microbial behavior, social physics and Origin of Life studies Acebrón et al. [2005]; García-Selfa et al. [2021]; Cooper et al. [2013]; Mimar et al. [2019]; COHEN [1977]; Rooney et al. [2006]; Xavier et al. [2020]; Blain and Szostak [2014]; Mimar et al. [2021].
@@
The KW18 formalism appears to offer an attractive path forward for the development of a working theory of semantic information. The difficulty, however, lies in the computational cost of applying the formalism to real world applications. Calculating the information theoretic measures require a precise elucidation of the state spaces, and the computation of the joint probability distributions between the agent and the environment. Unfortunately such state spaces are often of such high dimensionality that the formalism is computationally intractable. Thus if SIT is to live up to its promise, work must be done to find effective measures for the quantities required in its formalism Pemartín et al. [2024]. This includes best practices for definitions of viability, proper mechanisms for scrambling shared information in the intervened trajectories and, finally, practical methods of dimensional reduction (i.e. coarse-graining) for calculations of information theoretic metrics.
@@
Recently, an attempt to develop such effective versions of the KW18 formalism was presented in Sowinski et al. [2023]. The authors studied a forager model, a minimal model that captures attributes of living systems, namely exploration and resource consumption Bénichou and Redner [2014]; Bhat et al. [2017]; Bénichou et al. [2016]; Bhat and Redner [2022]. The viability was defined as the expected lifetime of the forager and the scrambling was done by noising a sensor that the forager used to detect resources. They applied and extended the work of KW18 and demonstrated how semantic information measures provide new insights into forager dynamics. More importantly, they provided a first cut at developing effective measures for application of the SIT formalism to real problems.
@@
An important result from Sowinski et al. [2023] was the identification of a plateau in the viability curve as a function of scrambling correlations between agent and environment. This plateau occurs at a viability threshold that is well below channel capacity. In other words a subset of correlations between agent and environment has no effect on the forager’s lifetime (the information was not semantic). Once the threshold is breached the ability for the forager to survive decreases monotonically with increased scrambling (noise in the sensor). Thus applying the ideas of SIT results in an identification of features in the viability curve that independent of the particular foraging strategy used. Whether such semantic saturation is a generic feature of living systems is far from settled, but the results make it clear that finding ways to apply the fundamental principle of SIT results in novel ways to analyze systems from an informational perspective.
@@
In this paper we report on the next step in efforts to operationalize SIT and explore effective versions of the theory. To carry out that task we explore the biophysically motivated problem of a control process (CP) acting on a base process (BP) to maintain overall system viability. We analyze a model of a network of coupled oscillators which tend towards synchronization (the base process) and an intervening agent (the control process) which uses the information on the network to limit synchronization. Our setting is loosely inspired by the neurodynamics of epileptic attacks Zheng et al. [2014]; Uhlhaas and Singer [2006]; Siegel et al. [2012]; Ghoshal et al. [2016] where brain pathologies occur due to large scale synchronization of neurons Penfield and Jasper [1954]. We study the synchronization dynamics through the well-known Kuramoto model of phase synchronization Kuramoto [1975], which while applicable to a wide range of settings, has also been used to describe complex brain wave activity: resting state functional connectivity Cabral et al. [2011], synchronization due to time-delayed signal transmission Petkoski and Jirsa [2019] and the effects of myelination Petkoski et al. [2023].
Additionally, in versions of the Kuramoto model with non-local coupling Abrams and Strogatz [2004] or symmetry-breaking Zhang and Motter [2021] there exist chimera states, where identically coupled oscillators coexist in synchronized and desynchronized states. Such states have been observed in both brain networks and the resuscitation of cardiac cells Omel’chenko [2018]. Thus desynchronization is important in many biological contexts. Consequently, the intervening agent here acts as preventing the system from reaching (a potentially) pathological synchronized state. Our goal in this paper, as in Sowinski et al. [2023], is to both explore the application of the KW18 formalism and use SIT to shed new light on a previously well-studied problem, phase-(de)synchronization.
@@
In what is to follow, in Sec. II we explicitly unpack the general steps required for the use of the SIT formalism. In Sec. III.1 we introduce the feedback-control model on the coupled oscillators. We demonstrate that introducing the agent leads to a stochastic version of the original Kuramoto model, with oscillators coupled to a heat-bath mediated by their degree $k_{i}$, and a deformed distribution of frequencies $g^{A}(\omega)$. Following this in Sec. III.2 we recast the model in information theoretic terms in which we show that the base process is a stochastic computation, and the control process is akin to analysing a communication channel between the agent and its environment. The viability function is defined in Sec. III.3 and in Sec. III.4 we analyze the semantic content of the correlations between agent/environment demonstrating that for a certain topology, there exists a semantic threshold, whereas for other connectivity patterns, there is no threshold, indicating that every bit of information in the communication channel is crucial to the viability of the system. Finally in Sec. IV we end with a discussion of the implications of our findings. Three Appendixes clarify technical aspects of our model and the approximations used.
@@ II Preliminaries
  • Dynamical System Decomposition. Construct a dynamical system $(S,\mathcal{D})$, where $S$ is the system state space and $\mathcal{D}:S\times T\rightarrow S$ is the dynamics on the state space for all times $t\in T$. The dynamics can be deterministic, stochastic, or a combination of the two; if they are continuous then $T\simeq\mathds{R}$ or some subinterval of the reals, and if they are discrete then $T\simeq\mathds{Z}$, or some subset of the integers. The system must be decomposable into an agent and environment, $S=A\times E$, meaning that any state of the system can be written $s=(a,e)$ where $a=(a_{1},a_{2},\dots)\in A$ and $e=(e_{1},e_{2},\dots)\in E$ are, respectively, the agent and environment dynamical degrees of freedom (DoF). We write $s(t)=(e(t),a(t))$ to represent the system evolution through state space, and refer to the specific form of $\mathcal{D}$ responsible for $a(t)$ as the agent’s behavior. $S$ needs to accommodate a (not necessarily unique) ground state for the agent which acts as an attractor for the system dynamics and represents a dead agent and an evolving agent-free environment.
@@
  • Coarse Grain and Quantify Correlations. Correlations between agent and environment DoF are manifest in the information architecture of the system. Information is quantified by a corpus of measures including mutual information, $I(A:E)$ Cover and Thomas [2006], cross entropy $H(A:E)$ Chen et al. [2022], and transfer entropy $\mathcal{T}_{E\rightarrow A}$ Schreiber [2000]. In practice these quantities may not be calculable or efficiently computable in full, so a bit of finesse is required at this step to find the right information measure to do the job Ibánez et al. [2024]; Tejero et al. [2024]. A combination of prior knowledge and intuition must be used to coarse grain the system into DoF that lend themselves to the desired analytical or computational techniques, while capturing the minimal dynamics of $S$ required to describe an agent attempting to survive in and co-evolve with its environment. In particular, the agent-free attractor must be preserved by the coarse-graining. For example, it would be intractable to describe a bacterium’s environment by listing the position and momentum of every atom surrounding it; a simpler description using a coarse-grained glucose density field might suffice. Describing the full biochemistry of the bacterium using every molecular DoF is intractable, but a coarse grained smaller set of DoF such as center of mass position, orientation, stored ATP density, and locomotive modality may not be. Note how this coarse-graining preserves the ground state: in the limit when the bacterium’s stored ATP supply vanishes it is effectively dead.
@@
  • Identify Viability Measure $\mathcal{V}$. Since the ground state remains an attractor for the system dynamics under coarse-graining, the distance to it endogenously defines the viability of an agent. Distance here can mean several things: If the attractor is stable, then an agent trajectory will eventually end up on it, i.e. the agent will die. In this case the expected lifetime of the agent is a proper measure of distance, and hence viability. The example of the bacterium falls in this case; the expected first crossing time to reach zero ATP supply is a viability measure. If the attractor is unstable, then the coarse-graining from the prior step must be able to identify an order parameter representing the efficacy of the agent at maintaining the agent-free out-of-equilibrium environment. The efficacy vanishes in an agent-free environment, and takes some non-zero value once the agent is introduced. The expected efficacy is then a proper measure of distance, and hence viability. The viability of the agent measured in either case is referred to as the actual viability.
@@
  • Scrambling Protocol and measuring semantic Content. The goal of SIT is to identify how the correlations and viability defined in the last two steps relate to one another. Specifically, to find which correlations matter to the agent. This is done by implementing a scrambling protocol on the correlations, and examining the resultant change in viability due to the impact the scrambled correlations have on the dynamics of the system. This was expounded on in KW18 using interventions to generate counterfactual histories, and requires a search over all possible partitions of the system degrees of freedom. Though rigorously defined, the procedure becomes unmanageable for systems with more than several hundred states. Rather than going through all partitions, the coarse-graining, if done correctly, identifies which coarse-grained DoF should be investigated and scrambled. In the example mentioned earlier, the bacterium’s orientation and whether it is tumbling or rolling are correlated to gradients and magnitudes of the glucose density. A scrambling protocol could be introduced at the level of the dynamics by intervening on the bacterium’s ability to sense either a magnitude or a gradient, or on the conditions that determine when the bacterium should switch it’s locomotive modality. The protocol should be tunable, in the sense that the actual viability is a limiting case. To contrast with the actual viability, the scrambled viability is measured as the protocol is tuned higher. The point at which the scrambled viability diverges from the actual viability is the semantic threshold Sowinski et al. [2023].
@@ III Kuramoto Model with Feedback-Control
Kuramoto’s famous coupled oscillator model has been well-studied since it’s introduction nearly half a century ago Kuramoto [1975]. From the synchronization of fireflies, cardiac pacemakers, laser, power transmission relay networks among others, it has proven a rich starting point for modelling systems in which synchronization is prominent Mirollo and Strogatz [1990]; Strogatz et al. [2005]; Moiseff and Copeland [2010]; Li and Wong [2017]. Above a critical coupling strength, an extensive fraction of oscillators achieve an equilibrium synchronized state, while the desynchronized states are non-equilibrium transients in the dynamics. As mentioned in the introduction, the phenomenon of desynchronization has biological relevance in terms of pathological states in the human brain and in cardiac arrhythmia. Inspired by this, we develop a feedback-control model of an agent acting on a population of coupled oscillators that attempts to prevent synchronization. We then apply the SIT framework to this idealized model and develop a new information theoretic perspective on coupled oscillators. In our formulation, the control process (an agent) seeks to drive the base process (the Kuramoto model) to a desired far-from-equilibrium set of states. Remaining far-from-equilibrium provides a definition of viability i.e. it motivates the teleological dynamics of the agent (the control process) seeking to keep the system from complete synchronization. Viewing the system through the lens of semantic information the agent can be interpreted as a repair process acting on a malfunctioning background.
@@ III.1.1 Environment Description
The environment is a population of $N$ oscillators $(\theta_{1}\dots\theta_{N})$, $\forall i,\theta_{i}\in(-\pi,\pi]$ placed on the nodes of an undirected graph $G$ with connections parameterized by the adjacency matrix ${{\boldsymbol{A}}}$ whose elements are $A_{ij}=1$ if nodes $i$ and $j$ are connected, and 00 otherwise. In what is to follow, we consider the fully-connected, Poisson and scale-free degree distributions generated via the configuration model Newman [2018]. The oscillators are coupled to one another with a coupling $\sigma$, which we take to be a function of time. Each oscillator has a natural frequency, $\omega_{i}$, drawn from a distribution $g(\omega)$. Without loss of generality we set $\langle\omega\rangle=0$. The variance, or more precisely the root mean square frequency $\omega_{\text{rms}}=\langle\omega^{2}\rangle^{1/2}$, is used to rescale time, $\tilde{t}=\omega_{\text{rms}}t$. Dropping the tildes, the dynamics of the dimensionless environment are the Kuramoto equations of motion, thus,
@@
with $g(\omega)$ now having vanishing mean and unit variance. Note, that the original variant of the model can be recovered by setting $A_{ij}=1$ (resulting in a fully-connected graph), and identifying $\sigma=K/N$, with $K$ being the unscaled coupling strength.